# Statistics of correlations in nonlinear recurrent neural networks

### v1

@@ 2.1 Collective fields
[…]
$Z_{N_{r}}(J)=\int d\phi\,d\eta\,d\rho\,e^{-\frac{i}{2}\eta_{ab}(f_{j}^{a}f_{j}^{b}-N\rho_{ab})}\,[\textrm{det}(2\pi M^{-1})]^{N/2}\,e^{-\frac{1}{2}(M^{-1})_{ab}\phi_{i}^{a}\phi_{i}^{b}}\,e^{J^{ab}_{i}f^{a}_{i}f^{b}_{i}}\,.$ (2.15)
@@
is independent of $\phi$. The equivalence with (2.13) follows from the fact that the integral over $\eta$ creates a delta function enforcing
$\rho_{ab}=\frac{1}{N}\sum_{j}f_{j}^{a}f_{j}^{b}$ (2.17)
These fields will play the role of collective coordinates describing the correlations of the RNN.
@@ 3 Large $N$ solution
$Z_{N_{r}}=\int d\phi\ d\eta d\rho\,\,e^{-\frac{i}{2}\eta_{ab}(\sum_{j}g_{ab}(\phi_{j})-N\rho_{ab})}\,[\textrm{det}(2\pi M^{-1})]^{N/2}\,e^{-\frac{1}{2}(G^{-1})_{ab}\phi_{i}^{a}\phi_{i}^{b}}e^{\sum_{j}J^{ab}_{j}(g_{ab}(\phi_{j})+\phi_{j}^{a}\phi_{j}^{b})}$ (3.2)
@@ 3.1 $1/N$ and source expansion
$\langle F(x)\rangle_{G}\equiv\int\frac{d^{N_{r}}x}{[\det(2\pi G)]^{1/2}}\,F(x)e^{-\frac{1}{2}(G^{-1})_{ab}x^{a}x^{b}}\,.$ (3.4)
@@
$Z_{N_{r}}(J)=\int d\eta d\rho\,\prod_{j}\langle e^{-\frac{i}{2}{\rm tr}(\eta g(x))+{\rm tr}\left(J_{j}(g(x)+xx)\right)}\,\rangle_{G}\,e^{\frac{i}{2}N\,{\rm tr}(\eta\rho)}[\textrm{det}(1+iM\eta)]^{-N/2}\,.$ (3.5)
@@
$\hskip-85.35826pt=\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}\,\textrm{exp}\!\left[\frac{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\left(J_{j}(g+xx)\right)\rangle_{G}}{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}}\right]$ (3.6)
$\langle\prod_{j}e^{-\tfrac{i}{2}{\rm tr}(\eta g)+{\rm tr}\left(J_{j}(g+xx)\right)}\rangle_{G}$ $=\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}\,\textrm{exp}\!\left[\sum_{j}\frac{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\,{\rm tr}\!\left(J_{j}(g+xx)\right)\rangle_{G}}{\langle e^{-\tfrac{i}{2}{\rm tr}(\eta g)}\rangle_{G}}\right]$ (3.7)
Here we will focus on calculating the partition function including second order contributions from the sources. We write Then
$Z_{N_{r}}(J)=\int d\eta d\rho\,e^{-NS_{eff}}\,.$ (3.8)
We will check self-consistently that the diagonal components of $\rho_{ab}$ and $G_{ab}$ are order $N^{0}$, while the off-diagonal components of $\rho_{ab}$ and $G_{ab}$, as well as all the variables $\eta_{ab}$ are order $1/N$. With this $N$-scaling, the leading terms at large $N$ are
@@
We have introduced the off-diagonal contribution $\sum_{a<b}\eta_{ab}^{2}\langle g_{ab}^{2}\rangle_{G}$ because it turns out to contribute at order $1/N^{2}$; this will have the same $N$-scaling as $\eta_{ab}\langle g_{ab}\rangle_{G}$, which is the leading contribution for $a\neq b$. In contrast, the leading contribution from $\eta_{ab}\langle g_{ab}\rangle_{G}$ for $a=b$, occurs at order $1/N$, while the corresponding quadratic term $\eta_{aa}^{2}\langle g_{aa}^{2}\rangle_{G}\sim 1/N^{2}$ and hence can be neglected. The same $N$-scaling arguments have been used for the terms we kept in the other terms of (3.9).
@@ 3.2 Effects of nonlinearities
In order to evaluate the $G$-expectation values in the effective action, we will approximate $G^{-1}$ in (3.4), taking into account that $G_{aa}\sim N^{0}$ while $G_{ab}\sim N^{-1}$ if $b\neq a$. Then
@@
In these and the following expressions, we are not summing over repeated replica indices.
With this approximation,
@@
For the off-diagonal component, the leadning nonvanishing contribution comes from expanding the Gaussian exponent to first order in $G_{a\neq b}$,
$\langle g_{a\neq b}\rangle_{G}\approx\int\frac{dx^{a}dx^{b}}{\sqrt{(2\pi G_{aa})(2\pi G_{bb})}}(f^{a}f^{b}-x^{a}x^{b})\left(\frac{G_{ab}}{G_{aa}G_{bb}}\right)e^{-\frac{(x^{a})^{2}}{2G_{aa}}-\frac{(x^{b})^{2}}{2G_{bb}}}\,.$ (3.13)
@@
$W(G_{aa})=\frac{1}{G_{aa}^{2}}\left(\langle g_{aa}^{2}\rangle_{W}+2\langle g_{aa}x^{a}x^{a}\rangle_{W}\right)\,.$ (3.19)
@@ 3.3 Saddle point description of neural correlation functions
which we recognize as a mean field equation. Such expressions have appeared before, e.g. in [24, 21]. It agrees with the self-consistent equation (2.2) found very recently in [18].
@@
Therefore, the 2-point function diverges as the coupling $\lambda\to 1$. This limit corresponds to the instability of the linear network, where the linear restoring term in (2.1) is overcome by the effects from fluctuations of th connectivity matrix. Understanding the network dynamics beyond this point requires including effects beyond the linear approximation. In our present formalism, we are including nonlinear effects exactly in the $1/N$ approximation; this divergence is resolved for activation functions that are bounded by $f(x)<x^{p}$ for $p<1$ at large $x$. To see this, Eq. (3.28) can be solved at large $G_{0}$ approximating the integral in $V(G_{0})$ by the regime where $f(x)\sim x^{p}$; the self-consistent solution then shows that the growth of $G_{0}$ is bounded by $G_{0}\lesssim(\textrm{const})\,\lambda^{2/(1-p)}$. See Sec 4.1 for more details.
@@
$\langle C_{ii}^{f}C_{ii}^{f}\rangle_{W}=\frac{1}{Z_{N_{r}}}\frac{\partial^{2}Z_{N_{r}}}{\partial J_{i}^{ab}\partial J_{i}^{ab}}\Big|_{J=0}\approx G_{0}^{2}V(G_{0})^{2}\approx\langle C_{ii}^{f}\rangle_{W}\,\langle C_{ii}^{f}\rangle_{W}\\ ,,$ (3.39)
@@
$\langle C_{ij}^{f}C_{ij}^{f}\rangle_{W}=\frac{1}{Z_{N_{r}}}\frac{\partial^{2}Z_{N_{r}}}{\partial J_{i}^{ab}\partial J_{j}^{ab}}\Big|_{J=0}\approx\frac{G_{0}^{2}}{N}\frac{(G_{0}-D)U(G_{0})^{2}V(G_{0})^{2}\left((D-G_{0})U(G_{0})^{2}+2G_{0}V(G_{0})\right)}{\left(G_{0}V(G_{0})-(G_{0}-D)U(G_{0})^{2}\right)^{2}}\,.$ (3.40)
@@
An important consequence of this property is that we now have a complete characterization of the probability distribution $P(C^{f})$ of the covariance matrix $C^{f}_{ij}$: its generating functional is given by a subset of the terms in (3.3),
@@
In other words, the statistics of $C^{f}$ is determined by the leading nonzero correlators (3.36), (3.40), plus large $N$ factorization.
@@ 3.4 Dimension of participation
3.5 Comparison to other works
On the other hand, Ref. [14] focused on the linear network. In this work a path-integral formulation is also used. In contrast to our work there are no replicas and a linear source term is used. Correlations up to order four are evaluated taking succesive derivatives with respect to the source and the mean values and variances of the elements of the covariance matrix are determined using a Wick decomposition of the temporal averages. This is possible in the linear case because the statistics for fixed couplings is Gaussian. This approach is not possible in the non-linear case, requiring the use of replicas.
The recent preprint [18] proposed an ansatz to solve the nonlinear network in terms of an effective coupling. This is motivated by mean field theory considerations and is verified numerically and using random matrix theory. This ansatz can be derived explicitly from our large $N$ treatment. From (3.50), the role of the effective coupling is played by
$\lambda_{eff}\equiv\lambda U(G_{0})\,.$ (3.52)
$U(G)=\int\frac{dx}{\sqrt{2\pi G}}\,f^{\prime}(x)\,e^{-\frac{x^{2}}{2G}}=\langle f^{\prime}(x)\rangle_{G}\,.$ (3.53)
The dimension of participation (3.50) becomes
$D_{PR}^{f}=N(1-\lambda^{2}\langle f^{\prime}(x)\rangle_{G}^{2})^{2}\,$, (3.54)
which coincides with Eq. (3.1) of [18] at vanishing frequency. Our result for $\langle C_{ij}C_{ij}\rangle_{W}$ is also consistent with theirs.
@@ 4.1 Power-law activations
The result for the participation dimension of the linear network has been obtained before in [19, 15]; the effective dimension vanishes as $\alpha^{2}\lambda^{2}\to 1$, corresponding to the onset of the instability of the linear network. On the other hand, for nonlinear networks, with $p<1$, this function does not vanish. The effective dimension is monotonically decreasing with $p$, attaining its maximum at $p=0$ (saturation regime), and vanishing for $p=1$ at the instability threshold.
@@ 4.2.1 Activation function with $p=0$
We find an excellent agreement between theory and simulation even for sizes of about a few hundred neurons. In Fig. 1 it is possible to see that the standard deviation of the different quantities decrease with network size. This is compatible with the $1/N$ expansion presented in Section 3.1. The fluctuations are controlled by higher order moments that fall with $N$ with $1/N^{2}$ or faster. In order to check this we performed simulations with sizes up to $N=800$ and evaluated the rate of decrease of the standard deviation of the average diagonal and off-diagonal terms of the output correlations. The results are shown on Fig. 2.
@@
In order to test the robustness of the results with respect to the simulation time we also performed simulations with a fixed number of steps $n_{t}$ (that corresponds to a simulation time $t_{sim}=n_{t}\delta$) without enforcing the condition $d<10^{-4}$. These results are shown in Fig. 3. We find thte results converge between $n_{t}=100$ and 200 iterations. This represents a simulation time between 10 and 20 times the characteristic time constant $\tau$.

### v2

@@ 2.2 Collective fields
Note that this is a symmetric matrix in the replica indices. Introducing an auxiliary field $\eta_{ab}$ that enforces this relation as a constraint ($\eta_{ab}$ is also symmetric in $(a,b)$), we arrive at the following representation for the path integral:
$Z_{N_{r}}(J)$ $=\int{\mathcal{D}}\phi\,{\mathcal{D}}\eta\,{\mathcal{D}}\rho\;\exp\!\left[-\frac{i}{2}\sum_{a,b=1}^{N_{r}}\eta_{ab}\Big(\sum_{j=1}^{N}f_{j}^{a}f_{j}^{b}-N\rho_{ab}\Big)\right]\,\big[\det(2\pi M^{-1})\big]^{N/2}$
$\qquad\times\exp\!\left[\sum_{i=1}^{N}\sum_{a,b=1}^{N_{r}}\left(-\frac{1}{2}(M^{-1})_{ab}\,\phi_{i}^{a}\phi_{i}^{b}+J_{i}^{ab}\,f_{i}^{a}f_{i}^{b}\right)\right].$ (2.17)
@@
is independent of $\phi$. The equivalence with (2.15) follows from the fact that the integral over $\eta$ creates a delta function enforcing (2.16). These fields will play the role of collective coordinates describing the correlations of the RNN.
@@ 3 Large $N$ solution
$Z_{N_{r}}(J)$ $=\int{\mathcal{D}}\phi\,{\mathcal{D}}\eta\,{\mathcal{D}}\rho\;\exp\!\left[-\frac{i}{2}\sum_{a,b=1}^{N_{r}}\eta_{ab}\left(\sum_{j=1}^{N}g_{ab}(\phi_{j})-N\rho_{ab}\right)\right]\,\big[\det(2\pi M^{-1})\big]^{N/2}$
$\qquad\times\exp\!\left[\sum_{j=1}^{N}\sum_{a,b=1}^{N_{r}}\left(-\frac{1}{2}(G^{-1})_{ab}\,\phi_{j}^{a}\phi_{j}^{b}+J_{j}^{ab}\,\big(g_{ab}(\phi_{j})+\phi_{j}^{a}\phi_{j}^{b}\big)\right)\right]$ (3.2)
@@ 3.1 $1/N$ and source expansion
$\langle F(x)\rangle_{G}\equiv\int\frac{d^{N_{r}}x}{[\det(2\pi G)]^{1/2}}\,F(x)e^{-\frac{1}{2}\sum_{a,b=1}^{N_{r}}(G^{-1})_{ab}x^{a}x^{b}}\,.$ (3.4)
@@
$Z_{N_{r}}(J)$ $=\int{\mathcal{D}}\eta\,{\mathcal{D}}\rho\;\exp\!\left[\frac{i}{2}N\,{\rm tr}(\eta\rho)\right]\,\big[\det(1+iM\eta)\big]^{-N/2}$
$\qquad\times\prod_{j=1}^{N}\left\langle\exp\!\left[-\frac{i}{2}{\rm tr}\!\big(\eta\,g(x)\big)+{\rm tr}\!\big(J_{j}\,(g(x)+xx)\big)\right]\right\rangle_{G}.$ (3.5)
@@
Our goal is to compute the partition function including second order contributions from the sources, and the dominant terms at large $N$. Let us write this as
$Z_{N_{r}}(J)=\int{\mathcal{D}}\eta\,{\mathcal{D}}\rho\,e^{-NS_{eff}}\,.$ (3.6)
The calculation of the “effective action” (in a statistical mechanics sense) is demonstrated in detail in Appendix C, and here we summarize the main steps. The first step is to write a cumulant expansion for the factor of (3.5) containing the sources. This can be done to all orders in powers of $\eta g$. The second step involves taking the large $N$ limit, and here comes a crucial simplification: it is possible to check self-consistently that the diagonal components of $\rho_{ab}$ and $G_{ab}$ are order $N^{0}$, while the off-diagonal components of $\rho_{ab}$ and $G_{ab}$, as well as all the variables $\eta_{ab}$ are order $1/N$. With this $N$-scaling, the leading terms at large $N$ are
@@
We have introduced the off-diagonal contribution $\sum_{a<b}\eta_{ab}^{2}\langle g_{ab}^{2}\rangle_{G}$ because it turns out to contribute at order $1/N^{2}$; this will have the same $N$-scaling as $\eta_{ab}\langle g_{ab}\rangle_{G}$, which is the leading contribution for $a\neq b$. In contrast, the leading contribution from $\eta_{ab}\langle g_{ab}\rangle_{G}$ for $a=b$, occurs at order $1/N$, while the corresponding quadratic term $\eta_{aa}^{2}\langle g_{aa}^{2}\rangle_{G}\sim 1/N^{2}$ and hence can be neglected. The same $N$-scaling arguments have been used for the remaining terms we kept in (3.7).
@@ 3.2 Effects of nonlinearities
In order to evaluate the $G$-expectation values in the effective action, we will approximate $G^{-1}$ in (3.3), taking into account that $G_{aa}\sim N^{0}$ while $G_{ab}\sim N^{-1}$ if $b\neq a$. [^6] At $J=0$, different replicas are identical but independent at leading order; the only source of replica coupling is the disorder average, which enters with the $1/N$ scaling of the weight covariance. Therefore the replica off-diagonal components are expected to vanish at leading order and to appear first as $O(1/N)$ corrections, a scaling that we will verify self-consistently at the saddle point. To leading order in the off-diagonal entries, this yields
@@
As before, sums over replica indices are always written explicitly. With this approximation,
@@
For the off-diagonal component the leading nonvanishing contribution comes from expanding the Gaussian weight to first order in $G_{ab}$. In particular, using 3.8 one has a cross-term $\exp\!\big[(G_{ab}/(G_{aa}G_{bb}))x^{a}x^{b}\big]\simeq 1+(G_{ab}/(G_{aa}G_{bb}))x^{a}x^{b}$. For odd activation functions (more generally, whenever $\langle f(x)\rangle_{G}=0$), the zeroth-order term vanishes and we obtain
$\langle g_{a\neq b}\rangle_{G}$ $\approx\int\frac{dx^{a}\,dx^{b}}{\sqrt{(2\pi G_{aa})(2\pi G_{bb})}}\,\Big(f^{a}f^{b}-x^{a}x^{b}\Big)\,\left(\frac{G_{ab}}{G_{aa}G_{bb}}\right)x^{a}x^{b}\,e^{-\frac{(x^{a})^{2}}{2G_{aa}}-\frac{(x^{b})^{2}}{2G_{bb}}}$ (3.11)
$=\frac{G_{ab}}{G_{aa}G_{bb}}\left(\langle xf(x)\rangle_{G_{aa}}\langle xf(x)\rangle_{G_{bb}}\right)-G_{ab}$ (3.12)
@@
Applying the same first-order approximation for the Gaussian weight, we obtain
$\langle g_{a\neq b}^{2}\rangle_{G}=\langle g_{a\neq b}^{2}\rangle_{0}+\frac{G_{ab}}{G_{aa}G_{bb}}\langle g_{a\neq b}^{2}x^{a}x^{b}\rangle_{0}+\mathcal{O}(G_{ab}^{2})\approx\langle g_{a\neq b}^{2}\rangle_{0},$ (3.15)
where $\langle\cdot\rangle_{0}$ denotes the decoupled Gaussian average. Expanding the integrand and using the factorization of the decoupled measure, we find
$\langle g_{a\neq b}^{2}\rangle_{G}\approx\langle f(x)^{2}\rangle_{G_{aa}}\langle f(x)^{2}\rangle_{G_{bb}}-2\langle xf(x)\rangle_{G_{aa}}\langle xf(x)\rangle_{G_{bb}}+\langle(x^{a})^{2}\rangle_{G_{aa}}\langle(x^{b})^{2}\rangle_{G_{bb}}\,.$ (3.16)
Finally, applying the definitions in Eqs. 3.10 and 3.13, we obtain
$\langle g_{a\neq b}^{2}\rangle_{G}\approx G_{aa}G_{bb}\left[V(G_{aa})V(G_{bb})-2U(G_{aa})U(G_{bb})+1\right]\,.$ (3.17)
@@
Substituting the explicit Gaussian averages computed above into Eq. (3.7), and using the decomposition for symmetric replica matrices ${\rm tr}(\eta A)=\sum_{a}\eta_{aa}A_{aa}+2\sum_{a<b}\eta_{ab}A_{ab}$, we rewrite each term in $S_{\rm eff}$. The combination appearing in the $\mathcal{O}(J)$ term becomes
$\langle g_{ab}(g_{ab}+x^{a}x^{b})\rangle_{G}=\langle g_{ab}^{2}\rangle_{G}+\langle g_{ab}\,x^{a}x^{b}\rangle_{G}\approx G_{aa}G_{bb}\Big(V(G_{aa})V(G_{bb})-U(G_{aa})U(G_{bb})\Big)$ (3.19)
@@
$W(G_{aa})=\frac{1}{G_{aa}^{2}}\left(\langle g_{aa}^{2}\rangle_{G}+2\langle g_{aa}x^{a}x^{a}\rangle_{G}\right)\,.$ (3.21)
@@ 3.3 Saddle point description of neural correlation functions
which we recognize as a mean field equation. Such expressions have appeared before, e.g. in [33, 31]. It has the same structure as the self-consistent equation (2.2) found very recently in [32]; we will compare with their results in Sec. 4.3.
@@
This result has been previously reported in [11]. Therefore, the 2-point function diverges as the coupling $\lambda\to 1$. This limit corresponds to the instability of the linear network, where the linear restoring term in (2.1) is overcome by the effects from fluctuations of the connectivity matrix. Understanding the network dynamics beyond this point requires including effects beyond the linear approximation. In our present formalism, we are including nonlinear effects exactly in the $1/N$ approximation; this divergence is resolved for activation functions that are bounded by $f(x)<x^{p}$ for $p<1$ at large $|x|$. To see this, Eq. (3.30) can be solved at large $G_{0}$ approximating the integral in $V(G_{0})$ by the regime where $f(x)\sim x^{p}$; the self-consistent solution then shows that the growth of $G_{0}$ is bounded by $G_{0}\lesssim(\textrm{const})\,\lambda^{2/(1-p)}$. See Sec 4.1 for more details.
@@
$\langle C_{ii}^{f}C_{ii}^{f}\rangle_{W}=\frac{1}{Z_{N_{r}}}\frac{\partial^{2}Z_{N_{r}}}{\partial J_{i}^{ab}\partial J_{i}^{ab}}\Big|_{J=0}\approx G_{0}^{2}V(G_{0})^{2}\approx\langle C_{ii}^{f}\rangle_{W}\,\langle C_{ii}^{f}\rangle_{W}$ (3.41)
@@
$\langle C_{ij}^{f}C_{ij}^{f}\rangle_{W}=\frac{1}{Z_{N_{r}}}\frac{\partial^{2}Z_{N_{r}}}{\partial J_{i}^{ab}\partial J_{j}^{ab}}\Big|_{J=0}\approx\frac{G_{0}^{2}}{N}\frac{(G_{0}-D)U(G_{0})^{2}V(G_{0})^{2}\left((D-G_{0})U(G_{0})^{2}+2G_{0}V(G_{0})\right)}{\left(G_{0}V(G_{0})-(G_{0}-D)U(G_{0})^{2}\right)^{2}}$ (3.42)
@@
An important consequence of this property is that we now have a complete characterization of the probability distribution $P(C^{f})$ of the covariance matrix $C^{f}_{ij}$. From Eq. (3.3) we retain only the terms that generate the leading cumulants relevant for $P(C^{f})$: the mean $\langle C^{f}_{ii}\rangle_{W}$, the local quadratic contribution, and the leading $O(1/N)$ variance of off-diagonal entries. The remaining terms in Eq. (3.3) generate cumulants that are not needed for $P(C^{f})$ at this order (or are fixed by large-$N$ factorization) and are therefore omitted in $Z_{C^{f}}(J)$. Finally, we use the self-consistent equation (3.30) to eliminate $\lambda^{2}$, which yields
@@
We conclude that in the large $N$ limit, the statistics of $C^{f}$ is determined by the leading nonzero correlators (3.38) and (3.42).
@@ 4.1 Power-law activations
The result for the participation dimension of the linear network has been obtained before in [12, 11, 19]; the effective dimension vanishes as $\alpha^{2}\lambda^{2}\to 1$, corresponding to the onset of the instability of the linear network. On the other hand, for nonlinear networks, with $p<1$, this function does not vanish. The effective dimension is monotonically decreasing with $p$, attaining its maximum at $p=0$ (saturation regime), and vanishing for $p=1$ at the instability threshold.
Figure 1: Comparison of analytical (dashed line) and numerical results (dots) for the output correlations. Left panel: average value of the diagonal correlation. Central panel: average of the square of the off-diagonal correlators. Right panel: participation ratio (participation dimension divided by network size). We use the transfer function of Eq.(4.1) with $\alpha=1$, $p=1/2$ and $D=1$; the network sizes are $N=50,100,200$. The error bars represent the standard deviation over 5 simulations with different realizations of the coupling matrix $W$ and noise $\xi$.
@@ 4.2.1 Activation function with $p=0$
We find an excellent agreement between theory and simulation even for sizes of about a few hundred neurons. In Fig. 2 it is possible to see that the standard deviation of the different quantities decrease with network size. This is compatible with the $1/N$ expansion presented in Section 3.1, since the fluctuations are controlled by higher order moments that fall with $N$ with $1/N^{2}$ or faster. A key point is that the observables we compare are strongly self-averaging: they involve averages over $\mathcal{O}(N)$ diagonal entries and $\mathcal{O}(N^{2})$ off-diagonal pairs (as well as averages over disorder realizations). As a consequence, the standard error decreases rapidly with N even before considering additional averaging over disorder realizations. In order to check this we performed simulations with sizes up to $N=800$ and evaluated the rate of decrease of the standard deviation of the average diagonal and off-diagonal terms of the output correlations. The results are shown on Fig. 3.
@@
In order to test the robustness of the results with respect to the simulation time we also performed simulations with a fixed number of steps $n_{t}$ (that corresponds to a simulation time $t_{sim}=n_{t}\delta$) without enforcing the condition $d<10^{-4}$. These results are shown in Fig. 4. We find the results converge between 100 and 200 iterations. This represents a simulation time between 10 and 20 times the characteristic time constant $\tau$.